OpenStudy (anonymous):

Which system of equations can be used to solve the following problem? Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold? Let c be the number of child tickets and a be the number of adult tickets.

OpenStudy (anonymous):

3x+5y=115 x+y=33

OpenStudy (anonymous):

y=33-x 3x+5(33-x)=115

OpenStudy (anonymous):

y=adult tickets x=child tickets

OpenStudy (anonymous):

here are the answer choices A. 3c+5a=115 a+c=35 B. 5c+3a=115 c+a=33 C. 3c+5a=33 a+c=115 D. 3a+5c=33 a+c=115

OpenStudy (anonymous):

@RyleeC

OpenStudy (anonymous):

Answer choice A is exactly the same thing I said in my first comment.

OpenStudy (anonymous):

oh i am sorry i am over thinking thanks